How much of potassium nitrate (in grams) would be needed to lower the melting point of 100 grams of water by 1 degree Celsius? Please show your work! Thank you.
Answer #1

∆T = 1°C

KNO3 = ( K+ & NO3- means i=2)

∆T = i* Kf * molality

Kf = 1.86°C/molality

1 = 2 * 1.86 * molality

Molality = 0.26882

Moles / mass of water ( in Kg) = 0.26882

Moles / ( 100 * 10^-3) = 0.26882

Moles = 0.026882

Mass of KNO3 = moles * molar mass

= 0.026882 * 101.1

= 2.718 grams

If you have any questions please comment

If you satisfied with the solution please rate it thanks